If a family decides to have children until their second boy is born what is the number of girls the family is expected to have?
Answer #1
here this follows negative binomial distribution with parameter r =number of boys required =2 and p=probabiltiy of a boy =0.5

expected number of total children =r/p=2/0.5 =4

since number of girls are 2 less than total number of children

therefore expected number of girls =E(x-2) =4-2 =2